The specific heat of aluminum is 0.902 j/g. oC. How much heat is absorbed by an aluminium pie tin with a mass of 473 g to raise its temperature from room temperature (23.00oC) to even temperature (375oF)?
1
Expert's answer
2020-10-20T08:52:20-0400

375oF =(375-32)*(5/9)oC

=190.5oC

Increase in temperature =190.5-23

=167.5

Required Heat= 0.902*167.5*473 J

=71463.205 J